OpenStudy (anonymous):

Find dy/dx for y=(sqrt x)(3x-1). I got the answer 9x-3/2sqrtx, but the answer is supposed to be 9x-1/2sqrtx. Can anybody help me understand what I did wrong?

myininaya (myininaya):

\[y=\sqrt{x}(3x-1)\]?

OpenStudy (anonymous):

yes

myininaya (myininaya):

Do you want to do the product rule and power rule? Or! Do you want to distribute and then do derivative junk?

OpenStudy (anonymous):

9x-1/2sqrtx is right

OpenStudy (anonymous):

i tried using the product and power rules.

myininaya (myininaya):

K \[(\sqrt{x})' =? , (3x-1)'=?\]

OpenStudy (anonymous):

1/2(x)^-1/2 , 3?

OpenStudy (anonymous):

\[\sqrt{x}(3x-1)=3x^{\frac{ 3 }{ 2 }}-x^{\frac{ 1 }{ 2 }}\]\[\frac{ d }{ dx }(3x^{\frac{ 3 }{ 2 }}-x^{\frac{ 1 }{ 2 }})=\frac{ 9 }{ 2 }x^{\frac{ 1 }{ 2 }}-\frac{ 1 }{ 2 }x^{\frac{ -1 }{ 2 }}\]

OpenStudy (anonymous):

Now rearrange the expression above so as to get what you miss.